OpenStudy (firejay5):

Solve and show and explain work x^2 + bx + c = 0

OpenStudy (anonymous):

Try applying the quadratic equation to this one.

OpenStudy (firejay5):

@LogicalApple

OpenStudy (anonymous):

\[x = \frac{ -b \pm \sqrt{b^{2} - 4ac} }{ 2a }\]

OpenStudy (firejay5):

@BubbaMurphy Do you have a different way to do this problem???

OpenStudy (anonymous):

LogicApple's way is definitely the way to do it.

OpenStudy (anonymous):

in this case, a=1

OpenStudy (firejay5):

Is that really how yo would solve it

OpenStudy (firejay5):

would A, B, and C = 1

OpenStudy (anonymous):

Based on the equation you supplied, a = 1, b = b, and c = c.

OpenStudy (anonymous):

They are coefficients of whatever quadratic you are talking about. Did you have a quadratic in mind?

OpenStudy (firejay5):

what you mean

OpenStudy (anonymous):

\[6x^{2} - 9x + 12 = 0\] \[-5x^{2} + x - 10 = 0\] \[x^2 + x - 1 = 0\] These are some examples of quadratics, all set equal to 0.

OpenStudy (firejay5):

Yea I know that

OpenStudy (anonymous):

Oh, ok... so what is your question?

OpenStudy (firejay5):

We have to solve x^2 + bx + c = 0 for bonus and I have no idea what to do

OpenStudy (anonymous):

I suggested using the quadratic equation. The only difference between x^2 + bx + c = 0 and the quadratics that I mentioned is that the coefficients in your equation are only known by their variable letter. But that is ok, you can still apply the quadratic formula to them. The solutions to x^2 + bx + c = 0 is: \[\frac{ -b \pm \sqrt{b^{2} - 4c} }{ 2 }\] This is the same thing I wrote before, except I let a = 1 because the coefficient of x^2 in your equation is 1.

OpenStudy (firejay5):

We didn't learn that equation you gave me

OpenStudy (anonymous):

What about completing the square ?

OpenStudy (firejay5):

so is that the answer

OpenStudy (anonymous):

Yes, but you would arrive at the same answer if you complete the square. Both techniques are valid for solving quadratic equations.

OpenStudy (firejay5):

I have to show my work, so how would I show it

OpenStudy (anonymous):

Well if you're using the quadratic equation you could say "applying the quadratic equation". But if you wanted to complete the square then: